what is the answer to this fraction problem please.
what is 1/3 of 9

Answer by jim_thompson5910(35256)

1/3 of 9 = (1/3)*(9) = (1*9)/3 = 9/3 = 3


So 1/3 of 9 is 3.


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim